### v1

# On Graph Odd Edge-Colorings and Odd Edge-Coverings

Abstract.
In this paper, we fully resolve two major conjectures on odd edge-colorings and odd edge-coverings of graphs, proposed by Petruševski and Škrekovski (European Journal of Combinatorics, 91:103225, 2021). The first conjecture states that for any loopless and connected graph $G$ with $\chi_{\text{odd}}^{\prime}(G)=4$, there exists an edge $e$ such that $G\backslash\{e\}$ is odd $3$-edge-colorable. The second conjecture states that any simple graph $G$ with $\chi_{\text{odd}}^{\prime}(G)=4$ admits an odd $3$-edge-covering in which only one edge receives two or three colors. In addition, we strongly confirm the second conjecture by demonstrating that there exists an odd $3$-edge-covering in which only one edge receives two colors.
1. Introduction
In this paper, we consider only undirected and loopless graphs, which may contain multiple edges unless specified otherwise as simple graphs. Throughout this paper, we employ the graph notation specified in [3].
For a graph $G$, an edge-coloring of $G$ is an assignment of colors from the set $[k]=\{1,2,\cdots,k\}$ to the edges of $G$, which provides an edge decomposition of $G$ into $k$ color classes. Similarly, an edge-covering of $G$ is an assignment of colors from $[k]$ to the edges of $G$, allowing for multiple colors per edge. Equivalently, we can denote an edge-coloring by a mapping $\varphi:E(G)\to[k]$. In the case of an edge-covering, we denote it by a set-valued mapping $\psi:E(G)\to 2^{[k]}\backslash\emptyset$. We will frequently talk about vertices with odd degrees or even degrees. For simplicity, we call these vertices odd vertices and even vertices, respectively. A graph is called an odd graph if every vertex is an odd vertex. We say $G$ is odd $k$-edge-colorable (respectively, odd $k$-edge-coverable) if there is an edge-coloring (respectively, an edge-covering) of $G$ by $k$ odd subgraphs, which we call an odd $k$-edge-coloring (respectively, an odd $k$-edge-covering). In other words, the coloring map or the covering map $\varphi$ is required to satisfy that for any $j=1,\cdots,k$, the edge set $E_{j}:=\{e\in E(G)\big{|}j\in\varphi(e)\}$ forms an odd graph. We also define $\chi_{\text{odd}}^{\prime}(G)$ (respectively, $\text{cov}_{\text{odd}}(G)$) to be the least integer $k_{0}$ such that $G$ is odd $k_{0}$-colorable (respectively, odd $k_{0}$-coverable). It is reasonable to call $\chi_{\text{odd}}^{\prime}(G)$ the odd chromatic index of the graph $G$. Alternatively, we understand $\chi_{\text{odd}}^{\prime}(G)$ as the least integer $k_{0}$ such that $G$ can be decomposed into $k_{0}$ (edge) disjoint odd subgraphs.
The topics of determining both $\chi_{\text{odd}}^{\prime}(G)$ and $\text{cov}_{\text{odd}}(G)$ for a graph $G$ have been very active in recent years, and we refer to the recent survey paper [6] for the history, references, and discussions on many related problems, ranging from theoretical problems on graphs and multigraphs to problems involving algorithms.
As an initiating result in this area, Pyber [8] proved the following theorem.
Theorem 1.1.
([8]) For any connected simple graph $G$, $\chi_{\text{odd}}^{\prime}(G)\leq 4$. If $G$ is a connected graph (not necessarily a simple graph) and it has even order, then $\chi_{\text{odd}}^{\prime}(G)\leq 3$.
It was also observed in [8] that the upper bound of $4$ in the previous result can be achieved by the wheel graph $W_{4}$. Lužar, Petruševski, and Škrekovski  [1] showed that connected loopless multigraphs are odd 5-edge-colorable, with one particular exception that is odd 6-edge-colorable. Before proceeding, we define a Shannon triangle of type $(2,2,2)$ as one of the graphs on three pairwise adjacent vertices, $v_{1},v_{2},v_{3}$, where any two of them are connected by multiple edges with even sizes. Similarly, we define a Shannon triangle of type $(2,2,1)$ for any graph on three pairwise adjacent vertices $v_{1},v_{2},v_{3}$, where both edge sets between the pairs $v_{1}v_{2}$ and $v_{1}v_{3}$ have even sizes, and the edge set between $v_{2}v_{3}$ has an odd size. It can be verified that if $S_{2,2,2}$ is a Shannon triangle of type $(2,2,2)$, then $\chi_{\text{odd}}^{\prime}(S_{2,2,2})=6$, and if $S_{2,2,1}$ is a Shannon triangle of type $(2,2,1)$, then $\chi_{\text{odd}}^{\prime}(S_{2,2,1})=5$. In 2017, Petruševski [4] proved that the above two families of graphs are the only exceptions; all other graphs are odd $4$-edge-colorable. Furthermore, Petruševski [4] also proved that when $\chi_{\text{odd}}^{\prime}(G)=4$, a color class can be reduced to a size of at most 2. In the end of the paper, the author raised a natural question regarding the existence of an edge-analogue of the second part of Theorem  1.1. That is, for a graph that is odd $4$-edge-colorable, is it possible to delete a particular edge (instead of a vertex), such that the resulting graph is odd $3$-edge-colorable? Additionally, the following conjecture was formally raised later in [6].
Conjecture 1.2.
(Conjecture 2.14 of [6]) For any connected graph which is not a Shannon triangle of type $(2,2,2)$ or type $(2,2,1)$, either $\chi_{\text{odd}}^{\prime}(G)\leq 3$, or there exists a special edge $e$ whose removal will make the graph odd $3$-edge-colorable.
In [4], an effort was already made towards the above conjecture. Note that an equivalent way to express the conjecture is that, assuming $\chi_{\text{odd}}^{\prime}(G)=4$, there exits an odd $4$-edge-coloring in which color $4$ is used only once. Another related result, present in [7], obtained a partial result by confirming the conclusion for the special class of subdivisions of odd graphs. Now we are ready to state the first result of this paper, which fully resolves the above conjecture positively.
Theorem 1.3.
For any connected graph $G$ that is not a Shannon graph of type $(2,2,2)$ or $(2,2,1)$, either $\chi_{\text{odd}}^{\prime}(G)\leq 3$ or $G$ admits an odd $4$-edge-coloring such that one of the four colors is used only once for some edge $e$. Clearly, in the latter case, $\chi_{\text{odd}}^{\prime}(G\backslash\{e\})=3$.
Now we move to graph coverability. Pyber also inquired in [8] whether three colors are sufficient to give an odd edge-covering for any simple graph $G$, a question later confirmed by Mátrai [2].
Theorem 1.4.
([2]) For any connected simple graph $G$, $\text{cov}_{\text{odd}}(G)\leq 3$.
This result was later partially generalized by Petruševski and Škrekovski [5] to multigraphs. Moreover, as a connection between $\chi_{\text{odd}}^{\prime}(G)$ and $\text{cov}_{\text{odd}}(G)$, as well as a tool to obtain more structural information about edge-colorings and edge-coverings, Petruševski and Škrekovski [5] defined the overlapping function. More precisely, for an edge-covering $\phi:E(G)\to 2^{[k]}\backslash\emptyset$, let $\text{ovl}(\phi)$ be the sum of sizes of all images under $\phi$ that are not singletons. The following are some examples of the values of the function $\text{ovl}(\phi)$:
$\text{ovl}(\phi)=\begin{cases}0,&\text{ if }\phi\text{ is an edge-coloring.}\\ 2,&\text{ if }\phi\text{ has exactly one edge receiving two colors, with others receiving one color.}\\ 3,&\text{ if }\phi\text{ has exactly one edge receiving three colors, with others receiving one color. }\\ 4,&\text{ if }\phi\text{ has exactly two edges receiving two colors each, withothers receiving one color.}\end{cases}$
Given a graph $G$ with $\text{cov}_{\text{odd}}(G)\leq 3$, define
(1.1) $\min\text{ovl}(G)=\min_{\phi}\text{ovl}(\phi),$
where the minimum is taken over all odd 3-edge-coverings $\phi$ of $G$. Immediately after the definitions, if $G$ is odd 3-edge-colorable, then $\min\text{ovl}(G)=0$. As the main result in [5], it was proved that except for a list of multigraphs with exactly $3$ vertices, for any other connected graph $G$, one has $\min\text{ovl}(G)\in\{0,2,3,4\}$. In the same paper, it was also shown that there do exist multigraphs $G$ for which $\min\text{ovl}(G)$ is at least $4$. So this leads to the following natural conjecture.
Conjecture 1.5.
(Conjecture 2.5 of [6]) There is no simple graph $G$ such that $\min\text{ovl}(G)=4$.
Our second result of this paper strongly confirms the above conjecture.
Theorem 1.6.
For any connected simple graph $G$, we have $\min\text{ovl}(G)\leq 2$.
Below, in Section 2, we start with a brief overview of notation and preliminary results. Subsequently, in Section 3, we will address some important special cases, obtaining certain partial results. We then proceed to prove Theorem 1.3 in Section 4 and Theorem 1.6 in Section 5.
2. Notation and Preliminary Results
In this section, we introduce some notation that we will frequently use in this work. Given a graph $G$, we define a bouquet $\mathcal{B}$ as the set of parallel edges on a pair of adjacent vertices. For any two adjacent vertices $v$ and $u$, let $\mathcal{B}_{vu}$ denote the edge set incident to $v$ and $u$. Given a loopless graph $G$, define a reduction graph $\text{red}(G)$ to be the spanning subgraph of $G$ where each bouquet $\mathcal{B}$ is reduced to size $1$ if it has an odd size, and to size $2$ if it has an even size. It was proved in [1] that the odd chromatic index of the reduction graph $\text{red}(G)$ is greater than or equal to that of $G$. It follows that for any loopless graph $G$, if $\text{red}(G)$ satisfies the conclusion of Theorem 1.3, then the same is true for $G$. Therefore, below we will only consider reduced graphs, that is, graphs satisfying $G=\text{red}(G)$.
A vertex $v$ is a cut vertex if $G-v$ is disconnected; otherwise, it is called an internal vertex. Let $v$ be a cut vertex of $G$, and let $C_{1},\cdots,C_{k}$ be the connected components of $G-v$. Then each $G[V(C_{i})\cup\{v\}]$, for $i=1,\cdots,k$, is called a $v$-lobe of $G$. A block of a graph $G$ is a maximal connected subgraph of $G$ that contains no cut vertices. For a block $B$ of a graph $G$, the collection of internal vertices in $B$ is denoted by $Int_{G}(B)$. If a block of $G$ contains at most one cut vertex of $G$, we say it is an end-block of $G$. For each vertex $v\in V(G)$, let $E_{G}(v)$ be the set of edges that are incident to $v$, and $N_{G}(v)$ be the set of neighbors of $v$ in $G$. In a reduced graph $G$, $N_{G}^{\prime}(v)$ denotes the set of neighbors of $v$ in a bouquet of size 1, and $N_{G}^{\prime\prime}(v)$ denotes the set of neighbors of $v$ in a bouquet of size 2. Let $T$ be an even-sized subset of $V(G)$. We define a spanning subgraph $H$ of $G$ to be a $T$-join if $d_{H}(v)$ is odd for all $v\in T$ and even for all $v\notin T$.
The following lemmas contain some long-known facts.
Lemma 2.1.
(Proposition 2.5 of  [4]) If $G$ is a connected graph that contains a bridge then $\chi_{\text{odd}}^{\prime}(G)\leq 3$.
Lemma 2.2.
([2] and [6]) Any forest $F$ is odd $2$-edge-colorable.
Moreover, for a fixed vertex $v$ in $F$, any local coloring of $E_{F}(v)$ that uses at most two colors extends to a 2-edge-coloring of $F$ that is odd away from $v$.
For the first statement above, we can start by considering any vertex $v\in F$. If $d_{F}(v)$ is odd, we can assign color $1$ to $E_{F}(v)$. If $d_{F}(v)$ is even, we can assign colors $1$ and $2$ to $E_{F}(v)$ arbitrarily, ensuring that both colors are used an odd number of times. Then, one can greedily extend the above coloring to an odd $2$-edge-coloring of $F$.
Lemma 2.3.
([9]) Let $G$ be a connected graph and let $T\subseteq V(G)$ be an even-sized subset. Then there exists a $T$-join of $G$ that can be either a forest or a coforest.
Let us remark that the second part of Theorem 1.1 follows directly from the above lemma.
Lemma 2.4.
(Proposition 2.6 of  [4]) Let $G$ be a connected graph and $v$ be an internal vertex. Let $T\subseteq V(G)$ be an even-sized set. Then there exists a $T$-join $H$ of $G$ which is a coforest, such that the following additional property holds for $v$ in $\hat{H}$, where $\hat{H}=G\backslash E(H)$.
  1. (a)
    If either $d_{G}(v)$ is odd and $v\in T$, or $d_{G}(v)$ is even and $v\notin T$, then $d_{\hat{H}}(v)=0$.
  2. (b)
    If either $d_{G}(v)$ is odd and $v\notin T$, or $d_{G}(v)$ is even and $v\in T$, then $d_{\hat{H}}(v)=1$.
Moreover, in Case (b), we can indicate a specific edge $e$ incident to $v$ such that $e\in E(\hat{H})$.
Recall that a block graph is a connected graph admitting no cut vertex. The following lemma is elementary and we omit its proof.
Lemma 2.5.
Let $G$ be a block graph with at least three vertices. Then for any three distinct vertices of $G$, there is a path which joins any two of them and contains the third vertex.
3. Some Special Cases
Aiming to prove our two main theorems, in this section, we present some preliminary results dealing with special cases. Note that some important partial results will already be obtained in this section. For example, Theorem 1.3 will be proved under the condition that the graph $G$ admits at least one cut vertex. We divide this section into two subsections for edge-colorings and edge-coverings, respectively. Certain arguments are based on the various coloring and covering schemes provided in [4] and [5], respectively.
3.1. Odd-Edge-Colorings
Let $S_{\ell}$ be a star graph with $\ell$ leaves. When $\ell=1$, we regard one vertex as a leaf and the other vertex as the center, instead of two leaves. We write $c(G)$ to denote the number of connected components of $G$. The following technical lemma assists in identifying an edge of interest in a fixed tree.
Lemma 3.1.
Let $T$ be a tree with $|V(T)|\geq 3$. Let $M$ be a subgraph of $T$, satisfying the following conditions.
  1. (1)
    each connected component of $M$ is either a copy of $S_{1}$ or a copy of $S_{2}$, whose leaves are also leaves of $T$.
  2. (2)
    the total number of leaves of all the connected components of $M$ is even.
  3. (3)
    $c(M)\geq 2$, and at least two connected components of $M$ are $S_{1}$’s.
Then there is an edge $e$ of $T\setminus E(M)$ such that $T\backslash\{e\}=L\cup R$ is a union of two subtrees, $L$ and $R$, and the number of leaves in both connected components $M\cap L$ and $M\cap R$ is odd.
Proof.
Assume for contradiction that $T$ is a tree with the smallest number of edges that admits some subgraph $M$ satisfying the three conditions, and the pair $(T,M)$ fails to satisfy the intended conclusion. By assumption, we can find two components of $M$ which are $S_{1}$’s. Let $u$ and $v$ denote their center vertices, respectively. Let $P_{uv}$ be the path in the tree $T$ joining vertices $u$ and $v$, and let $e$ be any edge in this path. Let us write $T\backslash\{e\}=L\cup R$, a union of two subtrees. By the assumption for $T$, the number of leaves in all the connected components of $M\cap L$ is even, and $M\cap L$ has at least one $S_{1}$. This implies that there are at least two $S_{1}$’s in $M\cap L$ and confirms condition (3). By the minimality of $T$, in $L$ we can find an edge $f$ such that $L\backslash\{f\}$ is a union of two subtrees $T_{1}\cup T_{2}$, each of which intersects $M$ with an odd number of leaves. Observe that exactly one of these two subtrees, say $T_{2}$, contains an end-vertex of $e$. Therefore, $T\backslash\{f\}$ is a union of two trees $T_{1}\cup(T_{2}\cup\{e\}\cup R)$, and the number of leaves of all the connected components of $T_{1}\cap M$ is odd, which is absurd. ∎
We define that a graph $H$ has a minimum feedback vertex set (abbreviated as $\min$ FVS) of size $1$ if $H$ contains a cycle and there exists a vertex $v$ such that $H-v$ is acyclic. We can refer to $v$ as a root of $H$. The following lemma is the same with Proposition 2.3 in [6].
Lemma 3.2.
Let $H$ be a graph with a $\min$ FVS of size $1$, with a root $v$. Then there exits an edge-coloring $\varphi:E(H)\to\{1,2\}$ such that for each $w\neq v$, both colors 1 and 2 are odd at $w$. Moreover, if $d_{H}(v)$ is odd, then color $1$ (respectively, color $2$) is odd (respectively, even) at $v$. If $d_{H}(v)$ is even, then either both color degrees of $v$ are even, or both are odd.
Lemma 3.3.
Let $H$ be a connected graph with a $\min$ FVS of size $1$, with a root $v$. Assume $\chi_{\text{odd}}^{\prime}(H)=3$, $d_{H}(v)$ is even, and $N_{H}^{\prime}(v)\geq 2$. Then there exists an edge $e$ of $H$ such that $H\backslash\{e\}$ is odd $2$-edge-colorable.
Proof.
Suppose $d_{H}(v)=2k$, for some integer $k\geq 1$. We obtain a new graph $H^{\prime}$ by blowing up the vertex $v$ into $2k$ distinct vertices, $v_{1},v_{2},\cdots,v_{2k}$. Therefore $H^{\prime}$ is a forest and the original edge set $E_{H}(v)$ becomes now $E(M)=\{e_{1},\cdots,e_{2k}\}$, where $M=H^{\prime}[\{v,v_{1},\cdots,v_{2k}\}]$. The components of $M$ are $S_{1}$’s and $S_{2}$’s, where there are at least two $S_{1}$’s since $N_{H}^{\prime}(v)\geq 2$. Clearly, by Lemma 2.2, $H^{\prime}$ admits an odd $2$-edge-coloring $\psi:E(H^{\prime})\to\{1,2\}$. The number of edges in $\{e_{1},e_{2},\cdots,e_{2k}\}$ that are assigned with color $1$ (or color $2$) is even; otherwise, $\psi$ induces an odd $2$-edge-coloring for $H$, contradicting our assumption that $\chi_{\text{odd}}^{\prime}(H)=3$. Now we consider the following cases.
Case 1. There is a component of $H^{\prime}$ containing an odd number of leaves of $M$.
Let $H_{1}$ be such a component, where there is an even number of edges of $M$ with one color, without loss of generality, color $1$, and an odd number of edges of $M$ with color $2$. Then, in $H^{\prime}-H_{1}$, there are also an even number of edges with color $1$ and an odd number of edges with color $2$. We can switch colors $1$ and $2$ in $H_{1}$ such that there are an even number of edges with color $2$ and an odd number of edges with color $1$. However, this leads to an odd $2$-edge-coloring for $H$, a contradiction with $\chi^{\prime}_{\text{odd}}(H)=3$.
Case 2. Every component of $H^{\prime}$ has an even number of leaves of $M$.
Since $N_{H}^{\prime}(v)\geq 2$, there is a component $H_{1}$ that has at least two $S_{1}$ stars in $M$. By Lemma 3.1, there is an edge $e$ of $H_{1}\backslash\{e_{1},\cdots,e_{2k}\}$ such that $H_{1}\backslash\{e\}$ has one component $H_{1}^{\prime}$ with an odd number of leaves in $M$. We can switch the edge colors in $H_{1}^{\prime}$, as in Case 1, and obtain an odd 2-edge-coloring of $H\backslash\{e\}$. Hence, $H\backslash\{e\}$ is odd 2-edge-colorable. ∎
The following two technical lemmas deal with graphs admitting cut vertices.
Lemma 3.4.
Suppose $G$ is connected with odd number of vertices and $\chi^{\prime}_{\text{odd}}(G)\leq 4$. Let $v$ be a cut vertex of $G$. The following are three cases to consider.
  1. Case 1.
    there exists a $v$-lobe $G_{1}$ such that $v(G_{1})$ is even.
  2. Case 2.
    there exists a $v$-lobe $G_{1}$ such that $d_{G_{1}}(v)$ is odd.
  3. Case 3.
    $v$ has even degree and $|N^{\prime}_{G}(v)|\geq 2$.
In each of the above cases, either $G$ admits an odd $3$-edge-coloring, or there exists an edge $e$ such that $G\backslash\{e\}$ admits an odd $3$-edge-coloring:
Proof.
Case 1. $G_{1}$ is a $v$-lobe such that $v(G_{1})$ is even. Let $G_{2}=G-(G_{1}-v)$, and note that $G_{2}$ is connected and $v(G_{2})$ is even. If $d_{G_{1}}(v)$ is even, we arbitrarily choose one edge $e\in E_{G_{1}}(v)$. Then we will prove that the graph $G\backslash\{e\}$ admits an odd $3$-edge-coloring. So now we assume that $d_{G_{1}}(v)$ is odd. By Lemma 2.4, we can take a spanning odd coforest $H_{1}$ for $G_{1}$, such that for the edge complement $\hat{H}_{1}=G_{1}\backslash E(H_{1})$, $d_{\hat{H}_{1}}(v)=0$. Similarly, we also take a spanning odd coforest $H_{2}$ of $G_{2}$, and therefore $\hat{H}_{2}=G_{2}\backslash E(H_{2})$ is a forest. Both $\hat{H}_{1}$ and $\hat{H}_{2}$ are odd $2$-edge-colorable. If $d_{G_{1}}(v)$ is odd, we have $d_{\hat{H}_{1}}(v)=0$. We assign $E(H_{1})$ with color $1$, and provide an odd $2$-edge-coloring to $E(\hat{H}_{1})$ with colors $2$ and $3$. We assign $E(H_{2})$ with color $2$, and assign color 1 (respectively, 3) to $E_{\hat{H}_{2}}(v)$ if $d_{\hat{H}_{2}}(v)$ is even (respectively, odd). In either case, we then extend the edge coloring to a $2$-edge-coloring for $E(\hat{H}_{2})$, which is odd-coloring at all vertices of $H_{2}$ away from $v$, via Lemma 2.2. Thus, the resulting coloring is indeed an odd $3$-edge-coloring for $G$.
Case 2. Suppose that there is a $v$-lobe $G_{1}$ such that $d_{G_{1}}(v)$ is odd and write $G_{2}=G-(G_{1}-v)$. By Case 1, we can assume that every $v$-lobe has odd order. In particular, $G_{1}$ has odd order. Let $H_{1}$ be a spanning odd coforest of $G_{1}-v$ and then $\hat{H}_{1}=G_{1}\backslash E(H_{1})$ is a graph with a $\min$ FVS of size $1$ with the root $v$. By Lemma 3.2, $\hat{H}_{1}$ admits a $2$-edge-coloring using colors $1$ and $2$, which is odd at all vertices $w\neq v$, $w\in V(G_{1})$, moreover, color $1$ is odd at $v$ and color $2$ is even at $v$. We assign color $3$ to $E(H_{1})$. If $d_{G_{2}}(v)$ is even, we choose an arbitrary edge $e\in E_{G_{2}}(v)$. By Lemma 2.1, $G_{2}\backslash\{e\}$ is connected. We only need to prove that $G\backslash\{e\}$ is odd 3-edge-colorable. In the following, we provide an odd 3-edge-coloring for $G$ when $d_{G_{2}}(v)$ is odd. We take a spanning coforest $H_{2}$ in $G_{2}-v$ and then $\hat{H}_{2}=G_{2}\backslash E(H_{1})$ is a graph with a $\min$ FVS of size $1$, with a root $v$. We assign $E(H_{2})$ with color $3$, and then use colors $1$ and $2$ to give $\hat{H}_{2}$ an $2$-edge-coloring, which is an odd coloring at every vertex $w\neq v$, $w\in V(G_{2})$, and color $1$ is even at $v$ and color $2$ is odd at $v$. This gives an odd $3$-edge-coloring for $G$.
Case 3. Considering both Case $1$ and Case $2$, we can assume that every $v$-lobe has an odd size, and $v$ has an even degree in every $v$-lobe. That is, each component of $G-v$ has an even size, and then we apply Lemma 2.3 to obtain an odd factor in each component. Thus, we denote the union of these odd factors as $F$ and assign color $1$ to $E(F)$. The graph $\hat{F}=G\backslash E(F)$ is either a forest, or a graph with a $\min$ FVS of size $1$, with a root $v$. If $\hat{F}$ is a forest, we can give it an odd $2$-edge-coloring, meaning that $G$ is odd $3$-edge-colorable. If $\hat{F}$ has a $\min$ FVS of size $1$, we can apply Lemma 3.3 to choose one edge $e$ of $\hat{F}$ such that $\hat{F}\backslash\{e\}$ is odd $2$-edge-colorable. This implies that $G\backslash\{e\}$ is odd $3$-edge-colorable, completing the proof. ∎
Lemma 3.5.
Suppose $G$ is connected with odd order and $\chi^{\prime}_{\text{odd}}(G)=4$. Assume $G$ admits a cut vertex $v$ with $N_{G}^{\prime}(v)=0$. Then there is an edge $e$ such that $G\backslash\{e\}$ admits an odd $3$-edge-coloring.
Proof.
We consider $G$ to be a minimal counter-example. Since $N_{G}^{\prime}(v)=0$, the degree of $v$ in every $v$-lobe is even. By Lemma 3.4, we can assume every $v$-lobe has odd order.
Let $G_{1}$ be an arbitrary $v$-lobe and $G_{2}=G-(G_{1}-v)$. If $G_{2}$ is of connectivity $1$, then the minimal choice of $G$ and Lemma 3.4 suggest that either $\chi_{\text{odd}}^{\prime}(G_{2})\leq 3$, or we can select an edge $e$ in $G_{2}$ such that $\chi_{\text{odd}}^{\prime}(G_{2}\backslash\{e\})=3$. In either case, we assume that color $1$ is assigned to some edge(s) incident to $v$ either in $G_{2}$ or in $G_{2}\backslash\{e\}$. Note that $v$ is an internal vertex of $G_{1}$. According to Lemma 2.4 (a), setting $T=G_{1}-v$, there exists a $T$-join $H$ of $G_{1}$, which is a coforest and $d_{\hat{H}}(v)=0$. We assign color 1 to $E(H)$ and obtain an odd $2$-edge-coloring of $\hat{H}$ with colors $2$ and $3$. Thus, we conclude that either $\chi_{\text{odd}}^{\prime}(G)\leq 3$, or we can find an edge (in $G_{2}$) such that $G\backslash\{e\}$ is odd 3-edge-colorable.
Next, we assume $G_{2}$ is a block graph. Let $u$ be a neighbor of $v$ in $G_{1}$ and $e\in\mathcal{B}_{vu}$. Write $G_{1}^{\prime}=G_{1}\backslash\{e\}$. We observe that $d_{G_{1}^{\prime}}(v)$ is odd and $d_{G_{2}}(v)$ is even. By Lemma 2.3, we can find an odd factor $H$ of $G_{1}^{\prime}-v$. Then, $\hat{H}=G_{1}^{\prime}\backslash E(H)$ is a graph with a $\min$ FVS of size $1$, for which $v$ is a root. Since $d_{\hat{H}}(v)$ is odd, Lemma 3.2 implies that $\hat{H}$ admits a $2$-edge-coloring odd away from $v$, where color $1$ is odd at $v$, while color $2$ is even at $v$.
Now, in $G_{2}$, we define the set $T=\{x\in G_{2}-v:d_{G}(x)\text{ is even}\}$. If $T$ is empty, then we assign colors $2$ and $3$ to each edge of the bouquet $\mathcal{B}_{wv}$, respectively, where $w$ is a neighbor of $v$ in $G_{2}$. All the other edges in $E(G_{2})$ are assigned color $1$. We can check that this yields an odd 3-edge-coloring for $G\backslash\{e\}$. If $T$ is not empty, according to Lemma 2.3, we can select a forest $F$ which is a $T$-join in $G_{2}$. If $d_{F}(v)=0$, we update $F$ to ensure that $v$ is incident to at least one edge of $F$. Specifically, let $u_{1}$ and $u_{2}$ be two vertices in $T$ such that there exists a path $P\subseteq F$ connecting them. By Lemma 2.5, there exists a path $Q$ connecting $u_{1}$ and $u_{2}$ passing through $v$. We then update $F$ by excluding $E(P)$, including $E(Q)$, and then deleting all the cycles (if any). It can be checked that $F$ remains to be a $T$-join, and now we confirm that $d_{F}(v)>0$. Let $\hat{F}=G_{2}\backslash E(F)$, where the only even vertex in $\hat{F}$ is $v$. Next, we apply an odd edge-coloring to the edges of $F$ using colors 2 and 3. By the previous argument, both colors $2$ and $3$ appear an odd number of times at $N_{F}(v)$. Therefore, we can assign the edges of $E(\hat{F})$ with color $1$. One can check that this procedure completes an odd $3$-edge-coloring of $G\backslash\{e\}$. ∎
We remark that the above two lemmas already imply Theorem 1.3 in the special case when the graph $G$ admits a cut vertex.
The following two lemmas are based on Corollary 3.3 and Lemma 3.4 in [4], respectively.
Lemma 3.6.
Let $G$ be a connected graph, and let $v$ and $w$ be odd vertices such that $w\in N_{G}^{\prime\prime}(v)$. If $v$ is an internal vertex of $G$ and $w$ is an internal vertex of $G-v$, then there exists an edge $f$ such that $G\backslash\{f\}$ admits an odd $3$-edge-coloring.
Proof.
Write $\mathcal{B}_{vw}=\{e,f\}$. Since $w$ is an internal vertex of $G-v$, by Lemma 2.4, there is a spanning odd coforest $H$ in $G-v$ such that $d_{\hat{H}}(w)=0$, where $\hat{H}=(G-v)\backslash E(H)$. We assign color $1$ to $E(H)$. Note that $d_{G}(v)$ is odd, and the graph $G\backslash E(H)$ has a $\min$ FVS of size $1$, with a root $v$. By Lemma 3.2, $G\backslash E(H)$ admits a $2$-edge-coloring using colors 2 and 3, which is an odd coloring away from $v$. Moreover, color $2$ is odd at $v$ and color $3$ is even at $v$. Clearly, we can assume that for the edge set $\mathcal{B}_{vw}=\{e,f\}$, edge $e$ receives color $2$ and edge $f$ receives color $3$. This induces a coloring of the graph $G\backslash\{f\}$, which is an odd $3$-edge-coloring. ∎
Lemma 3.7.
Let $G$ be a connected graph, and assume that every even vertex $u$ satisfies $N_{G}^{\prime}(u)=0$. Three vertices $v,w,q\in V(G)$ satisfy the following.
  1. (1)
    $v$ is an even internal vertex of $G$.
  2. (2)
    $q$ is an internal vertex of $G-v$.
  3. (3)
    $v$ and $q$ are nonadjacent, and $w$ is adjacent to both $v$ and $q$.
Then either $G$ admits an odd $3$-edge-coloring, or there is an edge $e$ such that $G\backslash\{e\}$ admits an odd $3$-edge-coloring.
Proof.
We discuss $4$ cases based on the parities of $d_{G}(w)$ and $d_{G}(q)$. We write $\mathcal{B}_{vw}=\{f_{1},f_{2}\}$ and specify one edge $e\in\mathcal{B}_{wq}$.
Case 1. $d_{G}(w)$ is odd and $d_{G}(q)$ is even. Since $q$ is an even internal vertex in $G-v$, according to Lemma 2.4, there is an odd spanning coforest $H$ in $G-v$, such that $E_{\hat{H}}(q)=\{e\}$, where $\hat{H}=(G-v)\backslash E(H)$. Note that $d_{\hat{H}}(w)$ is even, and $d_{\hat{H}}(w)\geq 2$ since $e$ is present. By Lemma 3.2, $G\backslash(E(H)\cup\{f_{1},e\})$ has a 2-edge-coloring $\phi$ such that $\phi$ is odd away from $v$, color $1$ is odd at $v$ and color $2$ is even at $v$. Since the degree of $w$ in $G\backslash(E(H)\cup\{f_{1},e\})$ is even, it has received both colors $1$ and $2$. Then we assign $\{e,f_{1}\}$ with color 2 and $E(H)$ with color 3. This procedure yields an odd 3-edge-coloring for $G$.
Case 2. $d_{G}(w)$ is even and $d_{G}(q)$ is odd. We write $\mathcal{B}_{wq}=\{e,e^{\prime}\}$. Note the graph $G\backslash\{e^{\prime}\}$ has a similar property with Case $1$, by exchanging the roles of $v$ and $q$. Therefore, a similar argument gives an odd $3$-edge-coloring for the graph $G\backslash\{e^{\prime}\}$.
Case 3. $d_{G}(w)$ is odd and $d_{G}(q)$ is odd. Choose $e^{\prime}\in E_{G}(q)\backslash E_{G}(w)$. Then $G\backslash\{e^{\prime}\}$ admits an odd $3$-edge-coloring as it satisfies condition of Case 1.
Case 4. $d_{G}(w)$ is even and $d_{G}(q)$ is even. We again write $\mathcal{B}_{wq}=\{e,e^{\prime}\}$. There exists a spanning odd coforest $H$ of $G-v$ satisfying $E_{\hat{H}}(q)=e$. Similar with Case 1, by Lemma 3.2, we can give a 2-edge-coloring $\phi$ with colors 1 and 2 for $G^{\prime}=G\backslash(E(H)\cup\{e,f_{1}\})$ such that $\phi$ is odd away from $v$, color $1$ is odd at $v$, and color $2$ is even at $v$. Note that $d_{G^{\prime}}(w)$ is odd, which implies that, via Lemma 2.2, we can ask $w$ to receive only one color in $\phi$. If $w$ receives color $1$, then we assign color $2$ for $f_{1}$, color $3$ for $E(H)$ and delete edge $e$, which is an odd $3$-edge-coloring for $G\backslash\{e\}$. If $w$ receives color $2$, then we assign color $2$ to both $f_{1}$ and $e$. It can be checked that this is an odd $3$-edge-coloring for $G$. ∎
We would like to address another very special case when $G$ is a complete graph with all adjacent vertices connected by double edges.
Lemma 3.8.
Suppose $G$ is a complete graph with order $12\ell+$ 1, $\ell\geq 2$, and each bouquet $\mathcal{B}_{vw}$ has size 2 for any $v,w\in V(G)$. Then, for any edge $e$, the graph $G\backslash\{e\}$ is odd $3$-edge-colorable.
Proof.
Let us specify two vertices $x$ and $y$ in $V(G)$, and let $H=G-\{x,y\}$, which is a copy of $K_{2\ell-1}$. Let $v$ be any vertex in $V(H)$. We can split $E(H)$ into a union of three edge sets (regarded as subgraphs): $H_{1}$, $H_{2}$, and an edge $e\in\mathcal{B}_{v_{1}v_{2}}$, where $v_{1}$ and $v_{2}$ are two fixed vertices in $H$. We require that the subgraph $H_{1}$ (respectively, $H_{2}$) has even degree at the vertex $v_{1}$ (respectively, $v_{2}$), and odd degree at all other vertices in $V(H)$. For example, we can take $H_{1}$ to be the union of a star $S_{2}$ centered at $v_{1}$ with two leaves $v_{2}$ and $v_{3}$, together with a perfect matching in $H-\{v_{1},v_{2},v_{3}\}$. Then $H_{2}=H\backslash(E(H_{1})\cup\{e\})$. We then define the graph $H_{1}^{\prime}$ to be the union of $H_{1}$ and all the bouquets $\mathcal{B}_{xw}$ for $w\in V(H)-v_{1}$ and one of the edges from the bouquet $\mathcal{B}_{xv_{1}}$. We can observe that $H_{1}^{\prime}$ is an odd graph. Similarly, we define $H_{2}^{\prime}$ to be a union of $H_{2}$ and all the bouquets $\mathcal{B}_{yw}$ for $w\in V(H)-\{v_{2}\}$, along with one of the edges from the bouquet $\mathcal{B}_{yv_{2}}$. The remaining graph $H_{3}=G\backslash(E(H_{1}^{\prime})\cup E(H_{2}^{\prime}))$ is the union of $\mathcal{B}_{xy}$ and one edge from $\mathcal{B}_{xv_{1}}$ and one edge from $\mathcal{B}_{yv_{2}}$, which is an odd subgraph. We assign color $1$ to $H_{3}$, color $2$ to $H_{1}^{\prime}$, and color $3$ to $H_{2}^{\prime}$. This is an odd $3$-edge-coloring for $G\backslash\{e\}$. ∎
The following lemma reveals an important structure where a particular even vertex $v$ satisfies $N^{\prime}_{G}(v)=0$ and is adjacent to every vertex of a path. There are two ways of a $2$-edge-coloring, both of which are odd at all vertices $w\neq v$. In the first, both colors $1$ and $2$ are odd at the vertex $v$, while in the other, both are even. This flexibility in coloring will be important later.
Lemma 3.9.
Let $G$ be a graph with $v(G)\geq 4$, and let $v$ be an even vertex such that $N_{G}(v)=N_{G}^{\prime\prime}(v)=V(G-v)$. Moreover, suppose $G-v$ is a simple path. Then we can obtain two distinct $2$-edge-colorings $\varphi_{1},\varphi_{2}:E(G)\to\{1,2\}$ as follows.
  1. (1)
    The first coloring $\varphi_{1}$ is an odd $2$-edge-coloring;
  2. (2)
    The second coloring $\varphi_{2}$ is a $2$-edge-coloring which is odd at all $w\neq v$, and both colors $1$ and $2$ are even at $v$.
Proof.
Let $\{e_{1},e_{2},\cdots,e_{2k}\}$ denote the edges that are incident to $v$, and let $P_{k-1}=G-v$ be the path. The edge set of $P_{k-1}$ is denoted by $\{f_{1},f_{2},\cdots,f_{k-1}\}$.
We define the first coloring $\varphi_{1}$ as follows. We assign color $1$ to the following edges.
$\{f_{1}\},\{f_{m}\big{|}\,m\text{ is even},2\leq m\leq k-1\},$
$\{e_{1},e_{2},e_{3}\},\text{ and }\{e_{2m+1},e_{2m+2}\big{|}\,m\text{ is even},2\leq m\leq k-1\}.$
The remaining edges are assigned color $2$.
We define the second coloring $\varphi_{2}$ as follows. We assign color $1$ to the following edges.
$\{f_{m}\big{|}\,m\text{ is odd},1\leq m\leq k-1\},$
$\{e_{1},e_{2}\},\,\text{ and }\{e_{2m+1},e_{2m+2}\big{|}\,m\text{is odd},1\leq m\leq k-1\}.$
The remaining edges are assigned color $2$. ∎
3.2. Odd-Edge-Coverings
Lemma 3.10.
Let $G$ be a connected simple graph with odd number of vertices. If $G$ admits a cut vertex $v$ such that $G-v$ has a connected component with odd order, then $\min\text{ovl}(G)\leq 2$.
Proof.
Let $C$ be a connected component of $G-v$ with $v(C)$ odd, and let $D=G-(V(C)\cup\{v\})$. By Lemma 2.3, there are odd spanning coforests $F_{1}$ and $F_{2}$ in $G[V(C)\cup\{v\}]$ and $G[V(D)\cup\{v\}]$ respectively. We also take them to have maximum sizes among all such odd spanning coforests. Next, we obtain $T_{1}=G[V(C)\cup\{v\}]\backslash E(F_{1})$ and $T_{2}=G[V(D)\cup\{v\}]\backslash E(F_{2})$. We assign colors $1$ and $2$ to the edges of $F_{1}$ and $F_{2}$, respectively.
Case 1. $d_{T_{1}\cup T_{2}}(v)\leq 1$.
We assume that $d_{T_{1}}(v)\leq 1$ and $d_{T_{2}}(v)=0$. Then, we proceed to provide an odd $2$-edge-coloring for $T_{1}$ using colors $2$ and $3$, ensuring that $E_{T_{1}}(v)$ does not receive color 2. Subsequently, we give an odd $2$-edge-coloring for $T_{2}$ using colors $1$ and $3$. As a result, we conclude that $\chi^{\prime}_{\text{odd}}(G)=3$, and thus, $\min\text{ovl}(G)=0$.
Case 2. $d_{T_{1}\cup T_{2}}(v)\geq 2$.
If $d_{G}(v)$ is odd, then $d_{T_{1}\cup T_{2}}(v)$ is also odd. Consequently, we can derive an odd $3$-edge-coloring for $T_{1}\cup T_{2}$, consistent with the coloring of $F_{1}$ and $F_{2}$ provided earlier. Initially, we ensure that all edges incident to $v$ are assigned color $3$. Subsequently, for the forest $T_{1}-v$, we utilize Lemma 2.2 to apply an odd $2$-edge coloring using colors $2$ and $3$. Similarly, for the forest $T_{2}-v$, we obtain an odd $2$-edge-coloring using colors $1$ and $3$. As a result, this yields an odd $3$-edge-coloring for the entire graph $G$, leading to $\min\text{ovl}(G)=0$.
If $d_{G}(v)$ is even, we define an odd $3$-edge-covering $\varphi$ such that $\text{ovl}(\varphi)=2$. Let $u\in F_{1}$ be a vertex adjacent to $v$. We assert that
Claim. $T_{1}\cup T_{2}\cup\{vu\}$ contains no cycle.
Indeed, if $T_{1}\cup T_{2}\cup\{vu\}$ contains a cycle, the cycle must be contained in $G[V(C)\cup\{v\}]$. Thus we can assume it has the form $vuw_{1}\cdots w_{t}$. Then, we can update $F_{1}$ by removing the edge $vu$ and adding the path $w_{1}\cdots w_{t}$, to obtain an odd coforest in $G[V(C)\cup\{v\}]$ which has larger size. This contradicts the maximum choice of $F_{1}$. The claim follows.
Then, by Lemma 2.2, we can obtain an odd $2$-edge-coloring for $T_{1}\cup T_{2}\cup\{vu\}$ with colors $3$ and $4$, such that all the edges incident to $v$ are colored with $3$ since $d_{T_{1}\cup T_{2}\cup\{vu\}}(v)$ is odd. Subsequently, we replace all edges colored with $4$ in $C$ (respectively, $D$) with color $2$ (respectively, color $1$). Thus, we obtain an odd $3$-edge-covering, where $vu$ is the only edge that receives two colors, namely, colors $1$ and $3$. Consequently, $\min\text{ovl}(G)\leq 2$. ∎
Below we rephrase Claim 2 and Claim 3 in the proof of Theorem 1 in  [5], only for simple graphs in our context.
Lemma 3.11.
Let $G$ be a connected simple graph with odd order and $v$ be an even vertex with $w\in N_{G}(v)$. We have the following:
  1. (a)
    If $d_{G}(w)$ is even and $w$ is an internal vertex of $G-v$, then $G$ admits an odd $3$-edge-coloring, which implies $\min\text{ovl}(G)=0$.
  2. (b)
    If $d_{G}(w)$ is odd, and there exists an internal vertex $u$ of $G-v$, such that $u\in N_{G-v}(w)$ and $u\notin N_{G}(v)$, then there is an odd $3$-edge-covering $\varphi$ of $G$, with $\text{ovl}(\varphi)\leq 2$.
Proof.
Let $d_{G}(v)=2k$.
(a) Since $w\in N_{G}(v)$ and $d_{G}(w)$ is even, we observe that $d_{G-v}(w)$ is odd. As $w$ is an internal vertex of $G-v$, we can find an odd factor $H$ of $G-v$, which is a coforest and contains the entire $E_{G-v}(w)$. Let $\hat{H}$ be the edge complement of $H$ in $G\backslash\{vw\}$. Clearly, $\hat{H}$ is a graph with a $\min$ FVS of size $1$, with $v$ as a root. Therefore, by Lemma 3.2, since $d_{\hat{H}}(v)$ is odd, there is a 2-edge-coloring $\varphi:E(\hat{H})\to\{1,2\}$ that is odd away from $v$, and for vertex $v$, color 1 is odd and color 2 is even. We then assign color $2$ to the edge $vw$ and color $3$ to $E(H)$. Thus, an odd $3$-edge-coloring for $G$ is obtained.
(b) We discuss two cases depending on the parity of $d_{G}(u)$.
Case $1$. $d_{G}(u)$ is even. In $G-v$, by Lemma 2.4(b), we can find an odd factor $H$, which is a coforest and contains $E_{G}(u)\backslash\{wu\}$. We assign color $3$ to $E(H)$ and let $\hat{H}$ be the edge complement of $H$ in $G\backslash\{vw,wu\}$. By Lemma 3.2, and observing that $d_{\hat{H}}(v)$ is odd, there is an edge-coloring $\varphi:E(\hat{H})\to\{1,2\}$ that is odd away from $v$, and for vertex $v$, color 1 is odd at $v$ and color 2 is even at $v$. If $w$ is an isolated vertex in $\hat{H}$, then we can assign color $2$ to the edge $vw$ and color $1$ to the edge $wu$. Otherwise, since $d_{\hat{H}}(w)$ is even, the edge set $E_{\hat{H}}(w)$ obtains both colors $1$ and $2$. We then assign color $2$ to both edges $vw$ and $wu$. This results in an odd $3$-edge-coloring for $G$, which implies $\min\text{ovl}(G)=0$.
Case $2$. $d_{G}(u)$ is odd. We find in $G-v$ an odd factor $H$, which is a coforest and contains $E_{G}(u)$. We assign color $3$ to $E(H)$ and let $\hat{H}$ be the edge complement of $H$ in $G\backslash\{vw,wu\}$. Again, by Lemma 3.2, and observing that $d_{\hat{H}}(v)$ is odd, there is an edge-coloring $\varphi:E(\hat{H})\to\{1,2\}$ that is odd away from $v$ with color $1$ being odd at $v$ and color $2$ being even at $v$. Note that $d_{\hat{H}}(w)$ is odd, so the edge set $E_{\hat{H}}(w)$ obtains exactly one of the colors $1$ or $2$. If it obtains color $1$, then we assign color $2$ to the edge $vw$, resulting in an odd $3$-edge-coloring $\varphi$, with $\text{ovl}(\varphi)=0$. If it obtains color $2$, then we assign color $2$ to both edges $vw$ and $wu$, resulting in the edge $wu$ receiving two colors in this case. In either case, the covering yields an odd $3$-edge-covering $\phi$ for $G$, with $\text{ovl}(\phi)\leq 2$. ∎
The next corollary is a rephrasing of Claim 7 in the proof of Theorem 1 in  [5].
Corollary 3.12.
Let $G$ be a connected simple graph with odd number of vertices. Suppose $v$ is an even internal vertex of $G$ and $G-v$ contains no cut vertices. Then $\min\text{ovl}(G)\leq 2$.
Proof.
Let $w$ be a neighbor of $v$, which is an internal vertex of $G-v$ by assumption. Lemma 3.11 (a) implies that $d_{G}(w)$ is odd, and Lemma 3.11 (b) implies that every neighbor $u$ of $w$ must also be a neighbor of $v$. By repeatedly applying Lemma 3.11 and noticing that $G-v$ is connected, we conclude that every vertex of $G-v$ has odd degree and is adjacent to $v$. Now we list the set of common neighbors of $v$ and $w$ as $\{q_{1},\cdots,q_{2k}\}$. We assign color $1$ to $wq_{j}$ and color $2$ to $vq_{j}$ for each $j=1,\cdots,2k$. Then we assign both colors $1$ and $2$ to the edge $vw$. Finally, note that the remaining edges form an odd graph, and we assign color $3$ to them. ∎
4. Proof of Theorem 1.3
This section is devoted to the proof of Theorem 1.3. The main strategy of the proof involves searching for different coloring schemes that use only one of the four colors once, given that $\chi_{\text{odd}}^{\prime}(G)=4$. For example, we have already seen that this strategy works for the special case when the graph admits a cut vertex, as discussed in the previous section. There are two additional ingredients in the proof. The first ingredient is the method of removing one edge from a certain tree or a graph with a $\min$ FVS of size $1$, as demonstrated in Lemma 3.3. This removal gives us some freedom to color the distinct components of the resulting forest separately. The second ingredient is a dichotomous approach to two coloring schemes for a specific graph, as demonstrated in Lemma 3.9. This flexibility is crucial for addressing the various cases that emerge in subsequent arguments.
Now we begin the proof by considering a connected graph $G$ which is not a Shannon triangle of type $(2,2,2)$ or type $(2,2,1)$. By Theorem 1.1, if $G$ has even order, then $\chi_{\text{odd}}^{\prime}(G)\leq 3$. If $v(G)=3$, it is also not hard to conclude. Therefore, below we can safely assume that $G$ has odd order and $v(G)\geq 5$. By comparing the parities, we observe that there exists at least one even vertex in $G$. Let $\mathcal{E}$ (respectively, $\mathcal{O}$) be the set of even (respectively, odd) vertices in $G$. Moreover, according to Lemmas 3.4 and 3.5, we can assume that $G$ is a block graph.
We then divide into two cases. In the first case, where there exists an even vertex $v$ satisfying $|N_{G}^{\prime}(v)|>0$, the proof will be concise since most of the groundwork has been covered in Section 3. In the second case, where every even vertex $v$ satisfies $|N_{G}^{\prime}(v)|=0$, we will follow some ideas from [4]. In particular, we will split the main argument into a sequence of six claims to organize various coloring schemes for different specific cases.
Case 1. There exists $v\in\mathcal{E}$ with $|N_{G}^{\prime}(v)|>0$.
The proof is similar to and simpler than the proof of Case 3 of Lemma 3.4. Thus, we will only outline the arguments here. We apply Lemma 2.3 to $G-v$ to obtain an odd spanning coforest $F$. Then, either $\hat{F}=G\backslash E(F)$ is a forest or it has a $\min$ FVS of size $1$, with a root $v$. We apply Lemma 3.3 to find the edge $e$ that satisfies the conclusion of Theorem 1.3.
Case 2. For every $v\in\mathcal{E}$, $|N_{G}^{\prime}(v)|=0$.
Claim 1. Let $v\in\mathcal{E}$ and $B$ be an end-block of $G-v$, then $\text{Int}_{G-v}(B)\subseteq N_{G}(v)=N_{G}^{\prime\prime}(v)$.
Proof.
By Lemma 3.7, we can see that every internal vertex of end-block $B$ is a neighbor of $v$. ∎
Claim 2. There exists a vertex $v\in\mathcal{E}$ such that $G-v$ is not a block graph.
Proof.
Suppose for a contradiction that for every vertex $v\in\mathcal{E}$ the graph $G-v$ is a block graph. By Claim 1, for each $v\in\mathcal{E}$, we have $N_{G}(v)=N_{G}^{\prime\prime}(v)=V(G)-v$.
If $|\mathcal{E}|=1$, we first observe that every vertex in $G-v$ has odd degree. So there must exist a pair of vertices $u,w$ such that $|\mathcal{B}_{uw}|=1$. Denote $e=uw$, and below we will show that removing $e$ from $E(G)$ results in a graph $G^{\prime}=G\backslash\{e\}$ that admits an odd $3$-edge-coloring. Since $G-v$ is a block graph, $G^{\prime}-v$ is connected and there is a path $P$ in $G^{\prime}-v$ connecting $u$ and $w$. We can see that $H=(G^{\prime}-v)\backslash E(P)$ is an odd graph. We assign color $3$ to $E(H)$. List the vertex set of $G-v-V(P)$ as $\{v_{1},v_{2},\cdots,v_{m}\}$ and write $\mathcal{B}_{vv_{i}}=\{e_{2i-1},e_{2i}\}$ for each $i=1,\cdots,m$. Then the edges of $\{e_{i}\big{|}\,1\leq i\leq 2m,i\text{ is odd}\}$ are assigned with color $1$ and $\{e_{i}\big{|}\,1\leq i\leq 2m,i\text{ is even}\}$ are assigned with color $2$.
Observe that by excluding isolated vertices, the remaining graph of $G^{\prime}\backslash E(H)$ has exactly the same structure as described in Lemma 3.9. If $m$ is even, we follow the first way in Lemma 3.9 to give it a $2$-edge-coloring. If $m$ is odd, we choose the second way. In either case, we eventually obtain an odd $3$-edge-coloring for $G^{\prime}=G\backslash\{e\}$.
For the case where $|\mathcal{E}|\geq 3$, according to Lemma 3.8, $G$ must have at least two odd vertices. Moreover, there must exist two adjacent odd vertices as $y$ and $z$, for which $|\mathcal{B}_{yz}|=1$. Let us select a triplet of vertices $v,w,x$ from $\mathcal{E}$. Now, we denote $\mathcal{B}_{vy}=\{e^{\prime},e^{\prime\prime}\}$ and $\mathcal{B}_{wz}=\{f^{\prime},f^{\prime\prime}\}$. We define $T=V(G)-\{v,w,x,y,z\}$ as the set containing an even number of vertices. Below we will define $H$ to be a $T$-join in $G-\{v,w,x\}$, which is also a coforest.
Specifically, let $p_{1},\cdots,p_{2\ell}$ and $o_{1},\cdots,o_{2k}$ represent the even and odd vertices in $T$, respectively. It is worth noting that $\ell$ and $k$ can be zero. When both $\ell$ and $k$ are zero, $H$ is an empty graph. Otherwise, by definition, the edge set $E(H)$ comprises the following parts:
  1. (1)
    $E(G[\mathcal{O}])\backslash\{yz\}$;
  2. (2)
    All the bouquets $\mathcal{B}_{p_{i}p_{i}^{\prime}}$, for $i,i^{\prime}=1,\cdots,2\ell$ and $i\neq i^{\prime}$, $\mathcal{B}_{p_{i}o_{j}}$, and $\mathcal{B}_{p_{i}z}$, for $i=1,\cdots,2\ell$, $j=1,\cdots,2k$;
  3. (3)
    one of the edges in each bouquet $\mathcal{B}_{p_{i}y}$, for $i=\cdots,2\ell$.
We then denote $\hat{H}=(G-\{v,w,x\})\backslash E(H)$ as the edge complement of $H$ in $G-\{v,w,x\}$. In fact, by definition, $\hat{H}$ is an odd star centered at $y$, with the vertices $z$, $p_{1},\cdots,p_{2\ell}$ being the leaves.
Consider the graph $G\backslash\{e^{\prime\prime}\}$. We assign color $1$ to $E(H)\cup\{e^{\prime},f^{\prime}\}\cup\mathcal{B}_{vz}$, color $2$ to $\mathcal{B}_{xz}\cup\{f^{\prime\prime}\}$, and color $3$ to $E(\hat{H})$. For each $p_{i}$, $i=1,\cdots,2\ell$, assign color $3$ to $\mathcal{B}_{p_{i}x}$ and also to $\mathcal{B}_{yx}$. For each $o_{j}$, $j=1,\cdots,2k$, the two edges in $\mathcal{B}_{o_{j}x}$ are assigned colors $2$ and $3$, respectively. It can be verified that so far there are an even number of colors 2 and 3 at $x$. Next, we assign the two edges in $\mathcal{B}_{vx}$ with colors $2$ and $3$, respectively, assign both edges in $\mathcal{B}_{wx}$ with color $2$, and assign all other still uncolored edges of $G$ with color $1$. One can verify that this yields an odd $3$-edge-coloring for $G\backslash\{e^{\prime\prime}\}$. ∎
Following Claim 2, we choose $v\in\mathcal{E}$ such that $G-v$ is not a block graph. For the next three claims, we fix an end-block $B$ of $G-v$ and denote by $s$ the unique cut vertex of $G-v$ that belongs to $V(B)$.
Claim 3. $\text{Int}_{G-v}(B)\subseteq\mathcal{O}$.
Proof.
Suppose $w$ is an even vertex in $\text{Int}_{G-v}(B)$. Let $B^{\prime}$ be another end-block of $G-v$, and $w^{\prime}\in\text{Int}_{G-v}(B^{\prime})$. Note that there is no edge incident to both $w$ and $w^{\prime}$, and $v$ is a common neighbor of $w$ and $w^{\prime}$. We apply Lemma 3.7 to the triple $\{w,v,w^{\prime}\}$ and conclude that $w^{\prime}$ is a cut vertex of $G-w$, which means that $G-\{w,w\}$ must be a disconnected graph. This has several strong implications. Firstly, $\text{Int}_{G-v}(B^{\prime})=\{w^{\prime}\}$. Then, together with Claim 1, it follows that $\text{Int}_{G-v}(B)=\{w\}$. Finally, $N_{G}(v)=N^{\prime\prime}_{G}(v)=\{w,w^{\prime}\}$. Recall that $s$ is the unique cut vertex of $G-v$ contained in $B$. Here we also denote by $s^{\prime}$ the unique cut vertex of $G-v$ in $B^{\prime}$ (note that it is possible that $s=s^{\prime}$). As a particular consequence of the above arguments, we see $v$ and $s$ (or $s^{\prime}$) are non-adjacent. Note that $|\mathcal{B}_{ws}|=2$ since $w$ is an even vertex.
In the graph $G-\{v,w,w^{\prime}\}$, we apply Lemma 2.3 to find an odd spanning coforest $H$, and write $\hat{H}=(G-\{v,w,w^{\prime}\})\backslash E(H)$. Choose one edge $e\in\mathcal{B}_{w^{\prime}s^{\prime}}$. Then we assign color $3$ to $E(H)$, and use colors $1$ and $2$ to give $E(\hat{H}+e)$ an odd $2$-edge-coloring such that the edge $e$ receives color $1$. Then, we assign color $3$ to both edges in $\mathcal{B}_{sw}$, assign color $1$ to both edges in $\mathcal{B}_{vw^{\prime}}$, and assign colors $1$ and $3$ to the two edges in $\mathcal{B}_{vw}$, respectively. If $|\mathcal{B}_{w^{\prime}s^{\prime}}|=1$, then we have already obtained an odd $3$-edge-coloring for $G$. Otherwise, let us denote by $e^{\prime}$ the other edge in $\mathcal{B}_{w^{\prime}s^{\prime}}$ other than $e$. Then, the graph $G\backslash\{e^{\prime}\}$ admits an odd $3$-edge-coloring. ∎
Claim 4. The graph $B-s$ is a simple graph.
Proof.
Suppose that we find vertices $w^{\prime}$ and $w^{\prime\prime}\in\text{Int}_{G-v}(B)$ such that $|\mathcal{B}_{w^{\prime}w^{\prime\prime}}|=2$. Note that by Claim 3, $w^{\prime},w^{\prime\prime}\in\mathcal{O}$. From Claim 1 we know that $v$ is adjacent to all vertices in $\text{Int}_{G-v}(B)$. And also $s$ and $v$ is connected in $G-\text{Int}_{G-v}(B)$. It follows that $G-w^{\prime}$ and $G-\{w^{\prime},w^{\prime\prime}\}$ are both connected. We then apply Lemma 3.6 to conclude the proof. ∎
Claim 5. $s\in\mathcal{O}$ and $\text{Int}_{G-v}(B)\cap N_{G}^{\prime\prime}(s)=\emptyset$.
Proof.
Suppose $s\in\mathcal{E}$. Lemma 3.7 implies that $\text{Int}_{G-v}(B)\subseteq N_{G}(s)=N_{G}^{\prime\prime}(s)$. Take any $w\in\text{Int}_{G-v}(B)$, then Claim 1 and Claim 4 imply $N_{G}^{\prime\prime}(w)=\{v,s\}$. Since $d_{G}(w)$ is odd, $N_{G}^{\prime}(w)$ consists of an odd number of internal vertices in $B$, listed as $\{u_{1},\cdots,u_{2\ell+1}\}$. For each $j=1,\cdots,2\ell+1$, choose one edge $e_{j}\in\mathcal{B}_{vu_{j}}$. Let $T$ be the set of all even vertices in the graph $G-v-(B-\{s\})$. Clearly, $T$ has even size and contains $s$. We apply Lemma 2.3 to find a forest $F$ which is also a $T$-join.
Let $\mathcal{B}_{ws}=\{e,e^{\prime}\}$. Now we consider the graph $G^{\prime}=G\backslash\{e^{\prime}\}$. With Lemma 2.2, we can assign color $1$ to all $E_{F}(s)$ and extend it to an odd $2$-edge-coloring to the whole $F$. We then assign color $1$ to $\{wu_{1},\cdots,wu_{2\ell+1}\}$, and assign color $2$ to $\{e_{1},\cdots,e_{2\ell+1}\}\cup\mathcal{B}_{vw}\cup\{e\}$. One can verify that the remaining edges of $G^{\prime}$ form an odd graph and assign color $3$ to it. Thus, an odd $3$-edge-coloring for $G^{\prime}=G\backslash\{e^{\prime}\}$ is obtained, and this shows $s\in\mathcal{O}$.
Now, Claim 3 and Lemma 3.6 imply that $\text{Int}_{G-v}(B)\cap N_{G}^{\prime\prime}(s)=\emptyset.$
We are now ready to make one final claim that will lead to the completion of our proof. This claim will address that there is at least one end-block whose order is at least $4$.
Claim 6. There is an end-block $B$ of $G-v$ with $v(B)\geq 4$.
Proof.
Firstly, an end-block $B$ cannot have exactly $1$ vertex by definition of a cut vertex. As stated in Claims 3, 4, and 5, an end-block $B$ also cannot contain exactly $3$ vertices since it is simple and every vertex in $B$ has an odd degree. Suppose now an end-block of $G-v$ contains exactly $2$ vertices, namely $u$ and $s$, where $u$ is an internal vertex and $s$ is a cut vertex of $G-v$. By Claims 3, 4, and 5, we know $d_{G}(u)=3$, $d_{G}(s)\geq 3$ is odd, and $|\mathcal{B}_{us}|=1$. Since $s$ is a cut vertex of $G-v$, it must be adjacent to some other neighbor other than $u$ and $v$ via an edge $e$. Below, we will show that $G^{\prime}=G\backslash\{e\}$ admits an odd $3$-edge-coloring.
Note that the vertex $s$ is not a cut vertex in $G^{\prime}$ and it is also not a cut vertex of $G^{\prime}-u$. In the graph $G^{\prime}-u$, since $d_{G^{\prime}-u}(s)$ is odd, Lemma 2.4(a) implies that there exists a spanning odd coforest $H$ such that $d_{\hat{H}}(s)=0$, where $\hat{H}=(G^{\prime}-u)\backslash E(H)$. Since $d_{G^{\prime}-u}(v)$ is positive and even, both $d_{H}(v)$ and $d_{\hat{H}}(v)$ are odd. We assign color $3$ to $E(H)$, and give an odd $2$-edge-coloring to $E(\hat{H})$ with colors $1$ and $2$. By Lemma 2.2, we can also ensure that all of $E_{\hat{H}}(v)$ receives color $1$. Finally, we assign color $1$ to $\mathcal{B}_{uv}\cup\{us\}$. This procedure results in an odd $3$-edge-coloring for $G^{\prime}=G\backslash\{e\}$. ∎
End of the proof of Theorem 1.3. Let $B$ be an end-block of $G-v$ containing at least $4$ vertices, and $s\in B$ be a cut vertex. Note that $d_{B}(w)\geq 3$, for $w\in Int_{G-v}(B)$. Let $G_{1}=G-v$. Hence, by applying Lemma 2.4, $G_{1}$ admits a spanning odd coforest $F$, and therefore $\hat{F}=G_{1}\backslash E(F)$ is a forest. We claim that $F$ contains all edges in $B$. In other words, the edge set of $\hat{F}$ does not intersect with the edge set of $B$. Suppose otherwise; then this intersection is again a forest and therefore has a leaf vertex in $B-s$ (or called a pendant vertex). It means there is a vertex $w\in\text{Int}_{G-v}(B)$ such that $d_{\hat{F}}(w)=1$. However, this implies that $d_{F}(w)$ is even and positive, which is absurd. The claim follows.
By the previous claims, $B-s$ is a simple connected graph with minimum degree at least 2. Then $B-s$ contains a cycle $C$. Let $F^{\prime}=F\backslash E(C)$ and note that $F^{\prime}$ is a spanning odd coforest of $G_{1}\backslash E(C)$. Let $e$ be an arbitrary edge of $C$, and let $P=C\backslash\{e\}$. Recall that $V(P)\subseteq N^{\prime\prime}_{G}(v)$. We will prove that $G\backslash\{e\}$ admits an odd 3-edge-coloring.
We assign color $3$ to $E(F^{\prime})$ and then apply Lemma 2.2 to give an odd $2$-edge-coloring to the forest $\hat{F}$, using colors $1$ and $2$. Consider any vertex $w\in(N_{G}^{\prime\prime}(v)-V(P))$. We extend the odd coloring to the edges of $\mathcal{B}_{vw}$. If the edge set $E_{G_{1}}(w)$ has received only color $3$, then the bouquet $\mathcal{B}_{vw}$ of two edges receives color $1$ and color $2$, respectively. Otherwise, if the edge set $E_{G_{1}}(w)$ has already received color $1$ (or $2$), we assign this color also to $\mathcal{B}_{vw}$. The remaining uncolored edges are the edges of $P$ and the edges connecting $P$ and $v$. These form a graph $H$, which has exactly the same structure as described in Lemma 3.9, leading to two possible cases.
  1. Case $1$.
    If both colors $1$ and $2$ appear an even number of times in the edge set $\bigcup_{w\in(N_{G}^{\prime\prime}(v)-V(P))}\mathcal{B}_{vw}$, we use the first method in Lemma 3.9 to assign colors to $E(H)$.
  2. Case $2$.
    If both colors $1$ and $2$ appear an odd number of times in the edge set $\bigcup_{w\in(N_{G}^{\prime\prime}(v)-V(P))}\mathcal{B}_{vw}$, we use the second method in Lemma 3.9 to assign colors to $E(H)$.
In either case, we can obtain an odd $3$-edge-coloring for the graph $G\backslash\{e\}$. ∎
5. Proof of Theorem 1.6
This section is devoted to the proof of Theorem 1.6. The proof, conducted by contradiction, begins with the consideration of a possible minimal counter-example and employs the technique of induction. The main inductive step is presented in the following proposition.
Proposition 5.1.
Let $G$ be a simple connected graph with $\min\text{ovl}(G)\geq 3$. Let $v\in V(G)$ be an even vertex.
  1. (a)
    If $v$ is a cut vertex of $G$, then one can find a connected component $C$ of $G-v$ such that the graph $G^{\prime}=G-C$ also satisfies that $\min\text{ovl}(G^{\prime})\geq 3$.
  2. (b)
    If $v$ is not a cut vertex, then $G-v$ is not a block graph. Furthermore, there is an end-block $B$ of $G-v$ containing a unique cut vertex $s$ of $G-v$ such that the graph $G^{\prime}=G-(B-s)$ satisfies that $\min\text{ovl}(G^{\prime})\geq 3$.
Proof.
(a) If any connected component of $G-v$ has an odd order, then Lemma 3.10 implies that $\min\text{ovl}(G)\leq 2$, which contradicts our assumption. Therefore, we can assume that all connected components of $G-v$ have an even order. There are two cases to consider.
Case 1. For every connected component $C$ of $G-v$, the degree $d_{G[V(C)\cup\{v\}]}(v)$ is odd.
In this case, since $d_{G}(v)$ is even, we conclude that there are an even number of components, namely, $C_{1},\cdots,C_{2k}$. In the graph $G[V(C_{1})\cup\{v\}]$, by Lemma 2.4(b), we can find a $V(C_{1})$-join $H_{1}$ which is a coforest and satisfies $d_{\hat{H}_{1}}(v)=1$, where $\hat{H}_{1}$ is the edge complement of $H_{1}$ in $G[V(C_{1})\cup\{v\}]$. We assign color $3$ to $E(H_{1})$. Clearly, $\hat{H}_{1}$ is a forest. We can then give an odd $2$-edge-coloring to $\hat{H}_{1}$ using colors $1$ and $2$, ensuring that $E_{\hat{H}_{1}}(v)$ receives color $1$.
For each of the remaining connected components $C_{j}$, $j=2,\cdots,2k$, we follow a similar process: we obtain a $V(C_{j})$-join $H_{j}$ for the graph $G[V(C_{j})\cup\{v\}]$ and assign color $1$ to it. Let $\hat{H}_{j}=G[V(C_{j})\cup\{v\}]\backslash E(H_{j})$ be the edge complement of $H_{j}$ in $G[V(C_{j})\cup\{v\}]$. We then give an odd $2$-edge-coloring to $\hat{H}_{j}$ using colors $2$ and $3$, such that $E_{\hat{H}_{j}}(v)$ obtains color 3.
It can be verified that the resulting coloring is an odd $3$-edge-coloring for the graph $G$. Consequently, $\min\text{ovl}(G)=0$, which contradicts the assumption.
Case 2. There is a connected component $C$ of $G-v$ such that $d_{G[V(C)\cup\{v\}]}(v)$ is even.
Let $G^{\prime}=G-C$ and we can see that $d_{G^{\prime}}(v)$ is even. Suppose, for contradiction, that $G^{\prime}$ has $\min\text{ovl}(G^{\prime})\leq 2$. Let $\varphi$ be an odd $3$-edge-covering of $G^{\prime}$ such that $\text{ovl}(\varphi)\leq 2$, and suppose color $3$ appears in $E_{G^{\prime}}(v)$ for $\varphi$.
Note that $v$ is an internal vertex for the induced subgraph $G[V(C)\cup\{v\}]$. Then, by Lemma 2.4(a), there exists a $V(C)$-join $H$ which is a coforest and contains the whole $E_{G[V(C)\cup\{v\}]}(v)$. In other words, we can define $\hat{H}$ to be the edge complement of $H$ in $G[V(C)\cup\{v\}]$, then $d_{\hat{H}}(v)=0$. We assign color $3$ to all edges in $E(H)$ and give an odd $2$-edge-coloring for $\hat{H}$ using colors $1$ and $2$. Clearly, this procedure extends $\varphi$ to an odd $3$-edge-covering of $G$, with overlapping at most $2$, which is a contradiction. This shows that $\min\text{ovl}(G^{\prime})\geq 3$.
(b) If $G-v$ contains no cut vertices, then Corollary 3.12 implies that $\min\text{ovl}(G)\leq 2$, which is absurd. So we can assume $G-v$ admits cut vertices. Let $B$ be an end-block of $G-v$, which contains a unique cut vertex $s$ of $G-v$. Note that every vertex $w\in B-s$ is an internal vertex of $G-v$. Therefore, Lemma 3.11 directly implies the following.
  1. (1)
    $v$ is adjacent to every vertex of $B-s$.
  2. (2)
    every vertex $w$ in $B-s$ satisfies that $d_{G-v}(w)$ is even.
We next claim that $s$ is adjacent to every vertex of $B-s$.
Suppose otherwise. Then we can find vertices $w\in B$ such that $s$ is not adjacent to $w$. Since $w$ has an odd degree, the neighbors of $w$ in $B-s$ can be listed as $\{q_{1},\cdots,q_{2\ell}\}$. Recall that $\mathcal{E}$ (respectively, $\mathcal{O}$) denotes the set of all even (respectively, odd) vertices of $G$. By comparing parities, we see $|\mathcal{O}|$ is even, implying that $|\mathcal{E}|$ is odd. We consider the even-sized vertex subset $T=\mathcal{E}-\{v\}$ in the graph $G-(\{v\}\cup(B-s))$. Applying Lemma 2.3, we obtain a $T$-join $H$, which is also a forest. We give an odd $2$-edge-coloring for $E(H)$ using colors 1 and 2. Then, we assign color $1$ to $wq_{j}$ and color $2$ to $vq_{j}$ for each $j=1,\cdots,2k$. For the special edge $vw$, we assign both colors 1 and 2. We check that all the uncolored edges of $G$ form an odd graph and then assign color $3$ to them. This defines an odd $3$-edge-covering $\varphi$, with $\text{ovl}(\varphi)=2$, which is a contradiction. Thus, the claim is established.
Now, in the graph $G^{\prime}=G-(B-s)$, we suppose, for contradiction, that it admits an odd $3$-edge-covering $\psi$ with $\text{ovl}(\psi)\leq 2$. Since the induced subgraph $B-s$ is an odd graph, it must have even order. Let $V(B-s)=\{v_{1},\cdots,v_{2k}\}$. In the covering $\psi$, if a color appears in both edge sets $E_{G^{\prime}}(v)$ and $E_{G^{\prime}}(s)$, let us say color $1$, then we can extend the covering to an odd $3$-edge-covering $\psi^{\prime}$ of $G$ with overlapping at most 2 by assigning color $1$ to all the uncolored edges. Otherwise, without loss of generality, we can assume that the edge set $E_{G^{\prime}}(s)$ contains color $1$ and does not contain color $3$, while $E_{G^{\prime}}(v)$ contains color $3$ and does not contain color $1$. Then, we assign color $3$ to all edges $vv_{j}$ and assign color $1$ to all edges $sv_{j}$, for $j=1,2,\cdots,2k$, and then assign color $2$ to all the edges in $B-s$. One can verify that this procedure extends $\psi$ to an odd $3$-edge-covering $\psi^{\prime}$ of $G$ with overlapping at most $2$. In either case, we have obtained a contradiction. ∎
Proof of Theorem 1.6.
With all the preparatory work completed as above, we proceed to prove the theorem by contradiction. To do this, we select a simple connected graph $G$ as our example, satisfying $\min\text{ovl}(G)\geq 3$ and having the minimum number of vertices. If $G$ contains at most $3$ vertices, then it contains at most $3$ edges, and clearly $\min\text{ovl}(G)=0$, which is absurd. Furthermore, since $v(G)=n$ is odd (as otherwise $\min\text{ovl}(G)=0$ by Theorem 1.1), we have $v(G)\geq 5$. By comparing parities, there exists at least one even vertex $v$. If $v$ is a cut vertex, we can apply Proposition 5.1 (a) to obtain an example $G^{\prime}$ with $\min\text{ovl}(G^{\prime})\geq 3$ with $|G^{\prime}|\leq|G|-2$, which contradicts our choice of $G$. If $G-v$ is connected, then we apply Proposition 5.1 (b) to obtain an example $G^{\prime}$ with $\min\text{ovl}(G^{\prime})\geq 3$ with $|G^{\prime}|\leq|G|-2$, which again contradicts our choice of $G$. ∎